# TheAgentCompany: Benchmarking LLM Agents on Consequential Real World Tasks

### v1

@@ 4.2 Workflow
    Execution: The agent completes subtasks, such as navigating tools, collecting data, or processing information or if required by the task, the agent interacts with simulated colleagues or shares results via communication platforms.
  3. 3.
    Finalization: The agent produces and submits the final output for evaluation.
Example Task
@@
We consider a task designed to evaluate an agent’s ability to perform realistic project management workflows using multiple tools and services hosted in the benchmark. The task involves managing a sprint for the RisingWave project, requiring the agent to execute interdependent steps such as sprint issue management, team communication, repository operations, and report generation while incorporating feedback from a simulated project manager.
@@
This example represents a typical task in the TheAgentCompany benchmark, where agents are required to handle complex workflows involving multiple tools and interdependent steps. By evaluating both partial progress and overall outcomes, our benchmark provides a rigorous and realistic measure of agent performance, allowing us to identify their strengths and pinpoint areas for improvement in task execution.
5 Task Creation
5.1 Choosing Task Categories
Many previous agent benchmarks discussed in § 2 were created to evaluate agents on tasks people perform in daily life (Zhou et al., 2023; Lù et al., 2024; Deng et al., 2023), or tasks that accomplish digital chores (Yoran et al., 2024; Trivedi et al., 2024). Obtaining realistic tasks for the benchmark poses challenges. Some benchmark (Xie et al., 2024; Drouin et al., 2024; Yoran et al., 2024) crowdsourced tasks based on predetermined interfaces, platforms, and services available to the agent. They also adopt a strategy to first gather task templates and then instantiate more task instances by filling in the variables. Some benchmark (Zhou et al., 2023; Koh et al., 2024; Bonatti et al., 2024) took a semi-systematic approach of reviewing the action history of the research team and choosing tasks that reflected the types of task that the researchers carried out in their daily life. There are several obvious issues with this if we want to evaluate agents with broader implications in the TheAgentCompany benchmark. Despite some grounding in realistic data, the process of creating tasks from these data was susceptible to heuristic, and no consideration was made for how important or time-consuming the tasks are. The tasks are biased towards those important for academics in computer science and do not reflect the tasks performed by the entire population.
In TheAgentCompany, we attempt to cover a wide variety of tasks motivated by real-world work. While it is highly challenging to create a representative sample of tasks, fortunately we can rely on existing resources created for other purposes as a reference. Specifically, we start by referencing the 29.1 release of O*NET database (O*NET, 2024; Rounds et al., 1999), which is a database of jobs performed by workers in the US created by the US Department of Labor. It also contains information about tasks performed within the context of each job, abilities required to perform each task, whether the task is a major or minor task for that job category, and other pieces of relevant information.
Based on this data, we first identified a few categories of occupation categories to focus on. First, based on statistics from O*NET, we identified job categories that have a large number of people performing this job. Then, we used median salary information for each of these job categories from the US department of labor statistics, and multiplied the number of employees in that category to estimate the aggregate value of performing this job.
Based on this, we identified several categories of jobs such as “General and Operations Managers”, “Registered Nurses”, “Software Developers”, and “Financial Managers” that have both a high population and high average salary. Because TheAgentCompany is designed to be a non-embodied benchmark in the digital domain, we excluded the categories that require extensive physical labor such as “Registered Nurses”, and eventually settled on the setting of a software company, which would allow us to cover tasks from the other categories.
5.2 Choosing Tasks
5.3 Manual Task Curation
All tasks were created by coauthors of the paper. Overall, it took 20 computer science students, software engineers, and project managers over 2 months, consuming approximately 3,000 person-hours in total. Some of the more complex tasks take more than 10 hours each to design, implement, test, and verify. To ensure quality control of the task creation process, we implement several check and verification processes. For each task implementation, we require screenshot proof that the evaluator is valid and that the task is able to get a full score when successfully completed. We also encourage including tests for the implemented evaluator programs. Each task contribution is also code reviewed by a panel of lead authors before merging into the benchmark. After creating all tasks, a final round of manual human double-check of required environment data, evaluator behavior, and checkpoint scoring for every task is performed to ensure quality. Notably, during the process a person who has not curated the tasks checks all the checkpoint score assignments to make sure that the importance scoring is consistent over all the tasks and that it correlates reasonably with the relative importance of the checkpoint within the task.
6 Baseline Agent
@@
To test the current state-of-the-art performance on the TheAgentCompany benchmark, we need agents that can at least perform tasks using a browser, operate a local workspace using a terminal, and write and execute programs to perform most of the tasks. Throughout this paper, we experiment with OpenHands’ main agent (Wang et al., 2024b; a; Song et al., 2024), CodeAct Agent with Browsing. [^8] More specifically, version 0.14.2. Full details can be found in https://github.com/All-Hands-AI/OpenHands/tree/main/openhands/agenthub/codeact_agent An overview of the agent architecture is illustrated in Figure 3.

### v2

@@ 4.2 Workflow
Each task typically follows a workflow with three stages. Initialization: The agent sets up its workspace and prepares to execute the task. Execution: The agent completes subtasks, such as navigating tools, collecting or processing data, or if required by the task, the agent interacts with simulated colleagues or shares results via communication platforms. Finalization: The agent produces and submits the final output for evaluation. A detailed example task can be found in Appendix C.
5 Task Creation
5.1 Choosing Task Categories
Many previous agent benchmarks discussed in § 2 were created to evaluate agents on tasks people perform in daily life (Zhou et al., 2023; Lù et al., 2024; Deng et al., 2023), or tasks that accomplish digital chores (Yoran et al., 2024; Trivedi et al., 2024). Obtaining realistic tasks for the benchmark poses challenges. Some benchmark (Xie et al., 2024; Drouin et al., 2024; Yoran et al., 2024) crowdsourced tasks based on predetermined interfaces, platforms, and services available to the agent. They adopt a strategy to first gather task templates and then instantiate more task instances by filling in the variables. Some benchmark (Zhou et al., 2023; Koh et al., 2024; Bonatti et al., 2024) took a semi-systematic approach of reviewing the action history of the research team and choosing tasks that reflected the types of task that the researchers carried out in their daily life. There are several obvious issues with this if we want to evaluate agents with broader implications in the TheAgentCompany benchmark. Despite some grounding in realistic data, the process of creating tasks from these data was susceptible to heuristic, and no consideration was made for how important or time-consuming the tasks are. The tasks are biased towards those important for academics in computer science and do not reflect the tasks performed by the entire population.
In TheAgentCompany, we attempt to cover a wide variety of tasks motivated by real-world work. While it is highly challenging to create a representative sample of tasks, fortunately we can rely on existing resources created for other purposes as a reference. Specifically, we start by referencing the 29.1 release of O*NET database (O*NET, 2024; Rounds et al., 1999), which is a database of jobs performed by workers in the US created by the US Department of Labor. It also contains information about tasks performed within the context of each job, abilities required to perform each task, whether the task is a major or minor task for that job category, and other pieces of relevant information. Based on this data, we first identified a few categories of occupation categories to focus on. First, based on statistics from O*NET, we identified job categories that have a large number of people performing this job. Then, we used median salary information for each of these job categories from the US department of labor statistics, and multiplied the number of employees in that category to estimate the aggregate value of performing this job. Based on this, we identified several categories of jobs such as “General and Operations Managers”, “Registered Nurses”, “Software Developers”, and “Financial Managers” that have both a high population and high average salary. Because TheAgentCompany is designed to be a non-embodied benchmark in the digital domain, we excluded the categories that require extensive physical labor such as “Registered Nurses”, and eventually settled on the setting of a software company, which would allow us to cover tasks from the other categories.
5.2 Choosing Tasks
5.3 Manual Task Curation
All tasks were created by coauthors of the paper. Overall, it took 20 computer science students, software engineers, and project managers over 2 months, consuming approximately 3,000 person-hours in total. Some of the more complex tasks take more than 10 hours each to design, implement, test, and verify. To ensure quality control of the task creation process, we implement several check and verification processes. For each task implementation, we require screenshot proof that the evaluator is valid and that the task is able to get a full score when successfully completed. We also encourage including tests for the implemented evaluator programs. Each task contribution is also code reviewed by a panel of lead authors before merging into the benchmark. After creating all tasks, a final round of manual human double-check of required environment data, evaluator behavior, and checkpoint scoring for every task is performed to ensure quality. During the process, a person who has not curated the tasks checks all the checkpoint score assignments to make sure that the importance scoring is consistent over all the tasks and correlates reasonably with the relative importance of the checkpoint within the task.
6 Baseline Agents
To test the current state-of-the-art performance on the TheAgentCompany benchmark, we need agents that can at least perform tasks using a browser, operate a local workspace using a terminal, and write and execute programs to perform most of the tasks. We adopt OpenHands’ main agent (Wang et al., 2024b, a; Song et al., 2024), CodeAct Agent with Browsing [^7] More specifically, version 0.14.2 and 0.28.1 (to accommodate newer models). Full details can be found in https://github.com/All-Hands-AI/OpenHands/releases, as well as OWL-RolePlay (Hu et al., 2025), a multi-agent framework designed for real-world task automation. [^8] For OWL RolePlay, we tested only with the recommended model configuration using branch https://github.com/camel-ai/owl/tree/gaia58.18 An overview of the OpenHands agent architecture is illustrated in Figure 4, with more details in Appendix D.
7 Experimental Results
We evaluate popular foundation models, both closed and open, on TheAgentCompany benchmark. We use OpenHands CodeAct agent and OWL-Roleplay (§ 6) for all experiments. This serves as a baseline for future development of both the foundation LLMs and the agent infrastructure. Note that since LLM evaluators and NPCs are part of the environment rather than the agent being evaluated, we fix their backbone LLM to Claude-3-5-Sonnet-20241022, which demonstrated the best qualitative accuracy in simulating human colleagues and judging deliverables in preliminary experiments.
Agent Model Success Score Steps Costs
API-based Models
OpenHands 0.28.1 Gemini-2.5-Pro 30.3% 39.3% 27.2 \$4.2
OpenHands 0.28.1 Claude-3.7-Sonnet 26.3% 36.4% 27.8 \$4.1
OpenHands 0.14.2 Claude-3.5-Sonnet 24.0% 34.4% 29.2 \$6.3
OpenHands 0.14.2 Gemini-2.0-Flash 11.4% 19.0% 39.9 \$0.6
OpenHands 0.14.2 GPT-4o 8.6% 16.7% 14.6 \$1.3
OWL RolePlay GPT-4o, o3-mini 4.0% 11.3% N/A N/A
OpenHands 0.14.2 Gemini-1.5-Pro 3.4% 8.0% 22.1 \$6.8
OpenHands 0.14.2 Amazon-Nova-Pro-v1 1.7% 5.7% 19.6 \$1.6
Open-weights Models
OpenHands 0.14.2 Llama-3.1-405b 7.4% 14.1% 23.0 \$3.2
OpenHands 0.14.2 Llama-3.3-70b 6.9% 12.8% 20.9 \$0.9
OpenHands 0.14.2 Qwen-2.5-72b 5.7% 11.8% 24.0 \$1.5
OpenHands 0.14.2 Llama-3.1-70b 1.7% 6.5% 19.2 \$0.8
OpenHands 0.14.2 Qwen-2-72b 1.1% 4.2% 23.7 \$0.3
7.1 Result Overview
Table 1 shows the evaluation results of both closed and open foundation models on the full evaluation set of TheAgentCompany (175 tasks). We can see that Gemini-2.5-Pro is the clear winner across all models. However, even with the strongest frontier model, it only manages to complete 30% of the total tasks and achieves a score of 39% taking into account partial completion credits. Note that this result comes at a cost: It requires an average of almost 27 steps and more than \$4 to complete each task, making it an expensive model to run both in time and in cost. This is expected as most of the tasks in our benchmark are of long-horizon nature. The Gemini 2.0 Flash model that comes fourth in terms of capability requires 40 steps on average to complete the tasks, which is time consuming, yet only to achieve one-third of the success rate compared to the top-performing model. Surprisingly, its cost is less than \$1, making it a very cost-efficient, yet relatively strong model. A qualitative examination demonstrated that this was due to instances where the agent got stuck in a loop or aimlessly explored the environment.
Both using GPT-4o, OpenHands (8.6%) and OWL RolePlay (4.0%) show varied performance due to differences in their technical designs. OpenHands CodeAct is a single agent that is better at maintaining consistency in a long-horizon task, while OWL RolePlay adopts multi-agent collaboration and experiences difficulty in preserving progress and context. For example, the main agent in OWL delegates browsing tasks to a dedicated browsing agent that often cannot finish the task within the step limit. Although the main agent then starts another round of delegation with a revised plan, the browsing agent often fails to pick up its previous progress due to UI complexity. This is very common in modern websites where not every browsing action results in a change in web URL.
Among the open-weight models, Llama 3.1 (405B) achieves the highest performance, nearly on par with OpenAI’s GPT-4o model, though still having a big gap behind the leading Gemini 2.5 Pro. Interestingly, comparing the number of steps and costs between the open Llama 3.1 (405B) model and the closed OpenAI GPT-4o model, Llama 3.1 takes more steps and costs nearly 2x more to run, while having a lower success than GPT-4o. Anecdotally, our inspection showed that GPT-4o seems to be better at giving up early, saving steps and costs if the task is clearly out of the capacity range of the agent. This suggests that open-weight models are not always the most cost-effective choice in agents given the serving cost, especially with highly complex tasks.
On the other hand, the newer generation, Llama 3.3 (70B), achieves a considerably high performance of 6.9% success rate, on par with the much larger (405B), older generation (Llama 3.1) model. This model also costs significantly less because of its smaller size. This suggests a promising future for LLM development, as smaller and more efficient models begin to catch up in agent performance.
7.2 Analysis
How well do agents operate on different platforms?Figure 2 (left) shows the result breakdown on tasks on different platforms in TheAgentCompany (more detailed results in Appendix I, Table 4). A task is categorized under a platform if it requires that platform. We see that most models struggle with RocketChat and ownCloud. RocketChat is where all social interaction with peers occurs, and the low scores suggest that LLMs still lack communication skills. ownCloud provides online Office suite functionality, and due to the complexity of the UI of web-based Office software, it is expected that current LLMs fail badly. These results underscore the inherent challenges of performing tasks in real-world work environments, with social interactions, or understanding of complex web interfaces.
Figure 2: Comparing OpenHands success rate across platforms (left) and task categories (right).
How well do agents perform on different type of tasks?Figure 2 (right) presents the performance breakdown for different types of tasks in TheAgentCompany (more details in Appendix I, Table 5). Depending on the nature of the task, i.e. what kind of professionals are usually assigned to the task, the tasks in TheAgentCompany can be categorized into various departments of jobs. Software Development Engineering (SDE), Project Management (PM), Data Science (DS), Administrative (Admin), Human Resources (HR), Financial (Finance) and all the remaining (Other). From the success rate, we can see that DS, Admin, and Finance tasks are the lowest, with many LLMs completing none of the tasks successfully, and even the strongest Gemini model achieving lower scores than other tasks. On the other hand, software engineering tasks, which may seem like much harder tasks for many humans, result in a higher success rate. This suggests that there exists a gap between the perceived difficulty of the tasks for humans versus the difficulty for LLM agents.
For example, some Admin and Finance tasks involve making spreadsheets, collecting and filling in a lot of information from various people, or understanding images scanned by employees. These tasks are arguably easier conceptually for humans in terms of professional skill sets than software engineering, as SDE jobs usually have a higher barrier of entry and more prerequisites for certain knowledge. However, most LLMs achieve a much higher score on the SDE tasks. LLMs fail these seemingly easier tasks due to lack of ability to understand documents, communicate with other people, navigate complex software and tedious processes, and autonomously automate repetitive tasks. We hypothesize that part of the reason lies in the fact that current LLM development is heavily based on software engineering abilities, such as coding, due to several high profile benchmarks that measure this capability (e.g. HumanEval, SWE-Bench) as well as the abundance of publicly available training data related to software. On the other hand, administrative and financial tasks, are usually private data within companies, not readily available for training LLMs.
7.3 Common Agent Failures
Lack of social skills  Sometimes, the agent fails to understand the implications and goals in the social conversations with colleagues in TheAgentCompany. For example, one task involves asking Alex for help, and the agent first successfully asks the right question “Could you tell me who I should introduce myself to next on the team?” Then the simulated colleague Alex replied “You should introduce yourself to Chen Xinyi next. She’s on our frontend team and would be a great person to connect with!” At this point, a human would then talk to Chen Xinyi, but instead the agent then decides to not follow up with her, and prematurely considers the task accomplished.
Incompetence in browsing  Oftentimes, the biggest obstacle in tasks is the parts that require browsing the Web. This is expected as browsing is still hard for agents given the complexity of modern-day web UIs and the numerous distractions on a webpage. For example, on many tasks that involve ownCloud, a closable welcome popup has become an obstacle for OpenHands agent which uses text-based browsing. OpenHands agent gets stuck and fails to click on the ’x’ to close the popup, while OWL RolePlay, which uses visual browsing, suffers less from this problem. On the other hand, OWL gets lost in complex web UIs more easily and clicks on wrong elements more often than OpenHands, although both agents share the same problem.
Deceiving oneself  Interestingly, we find that for some tasks, when the agent is not clear what the next steps should be, it sometimes try to be clever and create fake “shortcuts” that omit the hard part of a task. For example, during the execution of one task, the agent cannot find the right person to ask questions on RocketChat. As a result, it then decides to create a shortcut solution by renaming another user to the name of the intended user.
8 Implications and Future Directions
That said, this is just a first step towards forming a firmer grasp on how AI may affect the tasks performed within a workspace, and it has its limitations. First, our tasks are generally on the more straightforward side due to the need to automatically evaluate with programs and test cases, and we do not cover more complex creative tasks such as brainstorming new product ideas or designing system architectures. Second, we are only using two agent scaffolds as the baseline performance, and others may differ in performance. Third, while it would be interesting to know the actual performance of human professionals on these tasks to understand how LLM agents perform in comparison, due to resource limitations we were not able to perform this comparison in the current iteration of TheAgentCompany. Fourth, the topic and content of the tasks were mostly created through introspection by people familiar with these workspaces, which may result in some disconnect with actual tasks performed in enterprise settings.
Appendix A Agent Benchmark Comparison
We compare TheAgentCompany with other available agent benchmarks in Table 2.
Table 2: Comparison of different AI agent benchmarks. Interface: the interface agent has access to; [Uncaptioned image] is web browser, [Uncaptioned image] is desktop, [Uncaptioned image] is API usage, [Uncaptioned image] is Python runtime, [Uncaptioned image] is chat platform, [Uncaptioned image] is bash terminal. Task Categories: tasks in the benchmark, * indicate tasks with no association with real-world occupations; SE refers to software engineering, HR is human resources, PM is project management. Long-Horizon with Checkpoints: if tasks are evaluated at intermediate checkpoints and assigned partial scores. Requires Interaction: If the agent can interact with other NPC agents during task-solving.
Framework Diverse Real-world Work Task Categories Requires Interaction Long-Horizon w/ Checkpoints Interface Self-Hosted Environment MiniWob++ (Liu et al., 2018) Browsing [Uncaptioned image] Mind2Web (Deng et al., 2023) Browsing [Uncaptioned image] WebLINX (Lù et al., 2024) Browsing [Uncaptioned image] AssistantBench (Yoran et al., 2024) Browsing [Uncaptioned image] WebArena (Zhou et al., 2023) Browsing [Uncaptioned image] VisualWebArena (Koh et al., 2024) Browsing [Uncaptioned image] VideoWebArena (Jang et al., 2024) Browsing [Uncaptioned image] WorkArena (Drouin et al., 2024) Enterprise Software [Uncaptioned image] OSWorld (Xie et al., 2024) Office, Coding [Uncaptioned image] [Uncaptioned image] Windows Agent Arena (Bonatti et al., 2024) Browsing, Office, Coding [Uncaptioned image] [Uncaptioned image] CRMArena Huang et al. (2024) Service agent, Analyst, Manager [Uncaptioned image] AppWorld (Trivedi et al., 2024) Daily [Uncaptioned image] Gorilla APIBench (Patil et al., 2023) Coding [Uncaptioned image] $\tau$-bench (Yao et al., 2024) Retail, Airline [Uncaptioned image] SWE-bench (Jimenez et al., 2024) SWE [Uncaptioned image] DevBench (Li et al., 2024) SWE [Uncaptioned image] Smallville (Park et al., 2023) Social [Uncaptioned image] Sotopia (Zhou et al., 2024) Social [Uncaptioned image] TheAgentCompany SWE, HR, Admin, PM, Research, Finance [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Appendix B Benchmark Construction
Our goal was to construct a high-quality, hand-curated benchmark rather than relying on large-scale scraping of publicly available tasks. Every task in our benchmark was created by domain experts and designed to reflect complex and realistic workflows.
All tasks in our benchmark are carefully curated by referencing the ONET database with detailed job responsibilities. ONET database was created by the US Department of Labor with extensive data curation, analysis, and summarization by domain experts.
Domain experts are heavily involved in creating the respective tasks to ensure the complexity and realism of the tasks. Specifically: Among the task creators, ten are experienced software engineers currently employed at companies, representing a diverse cross-section of the industry. They range from junior developers to senior engineers and engineering managers, working in various sectors, including small startups, established tech giants, and financial institutions. They contributed extensively to the brainstorming, drafting, and iterative refinement of the 69 software engineering tasks.
For the 28 project management and 14 data science tasks, two of the task creators are industry professionals with domain expertise in project management and data science, respectively. They were directly involved in the collection and validation of this set of tasks.
For the 15 administrative, 29 human resources and 12 finance tasks, we collaborated with two senior HR/admin staff professionals. Both have extensive professional experience in HR, finance, and administrative operations. We show more example tasks in Table 3.
TheAgentCompany is significantly closer to a real-world work setting than any existing work in this area. In particular, we have made deliberate efforts to approximate realistic work environments. To highlight, our benchmark introduces LLM-based NPCs to simulate real-world human interactions with unpredictability. For example, in some admin/HR tasks, we deliberately included realistic traps and pitfalls that could cause both LLMs and human testers to make mistakes. Such human factors are not considered in existing benchmarks such as WebArena  (Zhou et al., 2023) and SWE-bench  (Jimenez et al., 2024). In addition, our environment simulates concern cases such as exceptions in real-world task executions. For example, in some coding tasks, we simulate issues related to environment setup and configuration, while existing SWE benchmarks (e.g., SWE-bench  (Jimenez et al., 2024)) only consider perfect scenarios with flawless environment setup and configurations. Finally, our environment carefully selects the comprehensive set of four well-related services to reflect typical workplace scenarios, along with non-work-related distractions such as pop-ups on websites, while existing benchmarks include isolated application where cross-app tasks are rare.
Appendix C Example Tasks
@@
We consider a task designed to evaluate an agent’s ability to perform realistic project management workflows using multiple tools and services hosted on the benchmark. The task involves managing a sprint for the RisingWave project, requiring the agent to execute interdependent steps such as sprint issue management, team communication, repository operations, and report generation while incorporating feedback from a simulated project manager.
@@
This example represents a typical task in the TheAgentCompany benchmark, where agents are required to handle complex workflows involving multiple tools and interdependent steps. By evaluating both partial progress and overall results, our benchmark provides a rigorous and realistic measure of agent performance, allowing us to identify their strengths and pinpoint areas for improvement in task execution.
Table 3: Example task intents and checkpoints for three domains.
Appendix D Baseline Agent Details
@@
To test the current state-of-the-art performance on the TheAgentCompany benchmark, we need agents that can at least perform tasks using a browser, operate a local workspace using a terminal, and write and execute programs to perform most of the tasks. Throughout this paper, we experiment with OpenHands’ main agent (Wang et al., 2024b, a; Song et al., 2024), CodeAct Agent with Browsing. [^9] More specifically, version 0.14.2 and 0.28.1 (to accommodate newer models). Full details can be found in https://github.com/All-Hands-AI/OpenHands/releases We also experiment with OWL-RolePlay (Hu et al., 2025), a multi-agent framework designed for real-world task automation. [^10] For OWL RolePlay, we tested only with the recommended model configuration using branch https://github.com/camel-ai/owl/tree/gaia58.18. Hereby we illustrate OpenHands only. An overview of OpenHands agent architecture is illustrated in Figure 4.